Laserwords Inc. is a book distributor that had been operating in its original facility since 1987. The increase in certification programs and continuing education requirements in several professions has contributed to an annual growth rate of 15% for Laserwords since 2012. Laserwords' original facility became obsolete by early 2017 because of the increased sales volume and the fact that Laserwords now carries CDs in addition to books.

On June 1, 2017, Laserwords contracted with Black Construction to have a new building constructed for $4,000,000 on land owned by Laserwords. The payments made by Laserwords to Black Construction are shown in the schedule below.

Date

Amount

July 30, 2017

$  900,000

January 30, 2018

 1,500,000

May 30, 2018

 1,600,000

Total payments

$4,000,000

Scenario 1: Construction was completed and the building was ready for occupancy on May 27, 2018. Laserwords had no new borrowings directly associated with the new building but had the following debt outstanding at May 31, 2018, the end of its fiscal year.

Scenario 2; The company borrowed an additional $1,000,000 at 10% interest to specifically fund the project.

  • 10%, 5-year note payable of $2,000,000, dated April 1, 2014, with interest payable annually on April 1.
  • 12%, 10-year bond issue of $3,000,000 sold at par on June 30, 2010, with interest payable annually on June 30.

The new building qualifies for interest capitalization. The effect of capitalizing the interest on the new building, compared with the effect of expensing the interest, is material.

Instructions

(a)  

Compute the weighted-average accumulated expenditures on Laserwords' new building during the capitalization period.

(b)  

Compute the avoidable interest on Laserwords' new building. (Round to one decimal place.)

(c)  Some interest cost of Laserwords Inc. is capitalized for the year ended May 31, 2018.

  • 1.Identify the items relating to interest costs that must be disclosed in Laserwords' financial statements.
  • 2.Compute the amount of each of the items that must be disclosed.
Answer #1

Scenario 1

a) Calculation of Weighted Average Accumulated Depreciation (Amounts in $)

Date Amount (A) Capitalization Period (B) Weighted Average Accumulated Expenditure (A*B)
July 30, 2017 900,000 10/12 months (Aug 17 to May 18) 750,000
Jan 30, 2018 1,500,000 4/12 months (Feb 18 to May 18 500,000
May 30, 2018 1,600,000 0/12 0
Total 4,000,000 1,250,000

Therefore the weighted-average accumulated expenditures on Laser words' new building during the capitalization period is $1,250,000.

b) Calculation of Weighted Average Interest rate (Amounts in $)

Loan Amount Actual Interest
Note Payable 2,000,000 (2,000,000*10%) = 200,000
Bonds 3,000,000 (3,000,000*12%) = 360,000
Total 5,000,000 560,000

Weighted Average Interest Rate = $560,000/$5,000,000 = 11.2%

Avoidable Interest = Weighted Average Accumulated Expenditure*Weighted Average Rate

= $1,250,000*11.2% = $140,000

c) Avoidable interest of $140,000 will be capitalized to the building and remaining interest will be charged to profit and loss account for the year ended May 31, 2018.

Total Actual Interest 560,000
Avoidable Interest 140,000
Total interest charged to P&L (560,000-140,000) 420,000

Scenario 2

a) The answer to part will remain same in this scenario also.

b) If company borrowed an additional special loan for the project then weighted average accumulated expenditure will be firstly used from special loan and remaining from borrowings outstanding. Calculation of avoidable interest in this case is shown as follows:-

Interest on Special loan = $1,000,000*10% = $100,000

Interest of expenditure used from beginning borrowing = ($1,250,000-$1,000,000)*11.2%

= $250,000*11.2% = $28,000

Avoidable Interest = $100,000+$28,000 = $128,000

c) Avoidable interest of $128,000 will be capitalized to the building and remaining interest will be charged to profit and loss account for the year ended May 31, 2018.

Total Actual Interest 560,000
Avoidable Interest 128,000
Total interest charged to P&L (560,000-128,000) 432,000